# Circuit-tuning: A Mechanistic Approach for Identifying Parameter Redundancy and Fine-tuning Neural Networks

### v1

@@ Abstract
The study of mechanistic interpretability aims to reverse-engineer a model to explain its behaviors. While recent studies have focused on the static mechanism of a certain behavior, the training dynamics inside a model remain to be explored. In this work, we develop an interpretable method for fine-tuning and reveal the mechanism behind learning. We first propose the concept of node redundancy as an extension of intrinsic dimension and explain the idea behind circuit discovery from a fresh view. Based on the theory, we propose circuit-tuning, a two-stage algorithm that iteratively performs circuit discovery to mask out irrelevant edges and updates the remaining parameters responsible for a specific task. Experiments show that our method not only improves performance on a wide range of tasks but is also scalable while preserving general capabilities. We visualize and analyze the circuits before, during, and after fine-tuning, providing new insights into the self-organization mechanism of a neural network in the learning process.
Machine Learning, ICML
@@ 1 Introduction
Benefiting from the high efficiency in computation and the scalable architecture of Transformer (Vaswani, 2017), large language models (LLMs) have demonstrated outstanding performance in a wide range of tasks. For real-world applications, fine-tuning aims to adapt a model to the downstream tasks. With the concept of intrinsic dimension (Li et al., 2018; Aghajanyan et al., 2020) that the parameters in a model are often redundant for solving a specific problem, parameter-efficient fine-tuning is proposed and has reduced computation to a great extent (Han et al., 2024). However, these techniques lack a mechanism fully clear to humans, and thus are less stable and may lead to catastrophic forgetting when applied to a new task (Wang et al., 2024). Besides, with the rapid development of LLMs, AI alignment has become an important issue in machine learning, highlighting the need for interpretability as well (Ji et al., 2023).
To discover the underlying mechanisms inside a model, (Elhage et al., 2021) provided a mathematical framework for transformer circuits, showing the potential of reverse-engineering a model under the guidance of mechanistic interpretability (Olah et al., 2020). Recent studies in this field have tried to isolate the circuit responsible for a single behavior (Wang et al., 2022), extract features using sparse autoencoders (Bricken et al., 2023), or apply a steering vector (Turner et al., 2023) to modify the model behaviors. While the mechanism of an existing model behavior is explored, few studies have paid attention to the mechanism in learning dynamics. Though several efforts have been made from various aspects (Olsson et al., 2022; Nanda et al., 2023), many issues are yet to be further explored in fine-tuning, e.g., how a new capability is acquired, how the model components interact and reorganize themselves, or how to intervene in the training process.
In this work, we develop an interpretable method for fine-tuning and reveal the mechanism behind learning. Our contributions are as follows:
  • We formulate different kinds of redundancy and propose the concept of node redundancy as an extension of the intrinsic dimension to the node level. Based on the theory, we provide an explanation for circuit discovery from the view of redundancy detection.
  • We propose circuit-tuning, an algorithm that iteratively finds out the components necessary for a task and performs fine-tuning in a mechanistic way. Our method not only performs well but is also scalable to various models and tasks while preserving general capabilities.
  • We provide a comprehensive analysis on training dynamics based on circuit-tuning and present some amazing findings for the self-organization inside a model, providing new insights and improvements for both fine-tuning and interpretability.
2 Background and Preliminaries
2.1 Features and Representations
Definition 2.1.
(Feature Vector) Given a representation space $\mathbb{V}^{D}$ with a set of bases $\mathbf{E}=\{\mathbf{e_{1},e_{2},...,e_{D}}\}$ where each basis $e_{i}$ corresponds to a neuron, the feature vector $\mathbf{v_{f}}$ of feature $f$ is a direction in $\mathbb{V}^{D}$, which is:
$\mathbf{v_{f}}=\sum_{i=1}^{m}c_{i}\mathbf{e_{i}}=c_{1}\mathbf{e_{1}}+c_{2}\mathbf{e_{2}}+...+c_{m}\mathbf{e_{m}}$
Remark 2.2.
The idea of Definition 2.1 follows the linear representation hypothesis that features are represented as vectors in a linear space, the evidence of which has been widely discussed in (Mikolov et al., 2013; Bricken et al., 2023; Park et al., 2024).
Assumption 2.3.
(Superposition) Given a neural network $\varphi$ and a representation space $\mathbb{V}^{D}$. If all features represented in $\varphi$ is $\mathcal{F}$, then we have $D<|\mathcal{F}|$, where $|\mathcal{F}|$ is the number of elements in $\mathcal{F}$.
Remark 2.4.
The idea comes from the superposition hypothesis discussed in (Olah et al., 2020; Elhage et al., 2022) which explains the phenomenon of polysemanticity that one neuron in a neural network often responds to multiple unrelated inputs. Since there are much more features than the dimensions in $\mathbb{V}^{D}$, a feature $f$ may not be aligned with a single basis. In fact, it is akin to the strategy of population coding (Pouget et al., 2000) in human brains that information is encoded by clusters of cells.
Considering the settings of Definition 2.1 and 2.3, we can have the conclusion: Given a task $T$ and an input $x$ with its features $\mathcal{F}_{x}=\{f_{1},f_{2},...,f_{t}\}\subseteq\mathcal{F}_{T}$. Suppose the feature vectors corresponding to the features in $x$ are $\mathbf{V_{x}}=\{\mathbf{v_{f_{1}}},\mathbf{v_{f_{2}}},...,\mathbf{v_{f_{t}}}\}$ and the activations of $\mathcal{F}_{x}$ are $A_{x}=\{a_{f_{1}},a_{f_{2}},...,a_{f_{t}}\}$, then the magnitude $a_{i}$ of $\mathbf{e_{i}}$ in the representation space $\mathbb{V}^{D}$ can be written as:
$a_{i}=\Big{(}\sum_{i=1}^{t}a_{f_{i}}\mathbf{v_{f_{i}}}\Big{)}\cdot\mathbf{e_{i}}$ (1)
This means the magnitude of dimension $i$ in $\mathbb{V}^{D}$ often originates from more than one features.
2.2 Computational Graph
If we view a model as a directed acyclic graph (DAG), then the nodes are terms in its forward pass (neurons, attention heads, etc.) and the edges are the interactions between the nodes. Note that the definitions of the node and edge do not necessarily follow the structure of the model. The shape of a node depends on the level of granularity when we inspect a model, and an edge can be a virtual connection between nodes far apart from each other.
2.3 Circuit Discovery

### v2

@@ Abstract
The study of mechanistic interpretability aims to reverse-engineer a model to explain its behaviors. While recent studies have focused on the static mechanism of a certain behavior, the learning dynamics inside a model remain to be explored. In this work, we develop an interpretable fine-tuning method for analyzing the mechanism behind learning. We first introduce the concept of node-level intrinsic dimensionality to describe the learning process of a model in a computational graph. Based on our theory, we propose circuit-tuning, a two-stage algorithm that iteratively builds the minimal subgraph for a specific task and updates the key parameters in a heuristic way. Experimental results confirm the existence of the intrinsic dimensionality at the node level and demonstrate the effectiveness of our method for transparent and interpretable fine-tuning. We visualize and analyze the circuits before, during, and after fine-tuning, providing new insights into the self-organization mechanism of a neural network in the learning process.
@@ 1 Introduction
Transformer-based large language models (LLMs) have demonstrated outstanding performance in a wide range of tasks [61]. However, an LLM is often treated as a "black box” because of its complex inner mechanisms, which brings a lot of issues about AI saftey and reliability, highlighting the need for interpretability [24]. Mechanistic interpretability aims to discover the underlying mechanisms inside a model so as to provide better control and improved design of it [53], showing the potential of reverse-engineering a model. Recent studies in this field include analyzing the circuit responsible for a single behavior [41, 62], extracting features via sparse dictionary learning [5, 57], applying a steering vector [60] to modify the model behaviors, etc.
Despite the success of the above methods, they are limited to post-hoc analyses of a trained model. For example, Wang et al. [62] and Hanna et al. [18] studied the interactions of the attention heads / MLPs and discovered the circuits for indirect object identification and mathematics in GPT-2-small. This kind of static analysis of model behaviors during inference fails to explain the learning dynamics of a model, i.e., how a model acquire an ability, or generalize to various scenarios, which is of great importance in mechanistic interpretability [53]. Recently, Nanda et al. [38], Olsson et al. [42] studied the phase transition during training. Prakash et al. [46], Lee et al. [28], Jain et al. [23], Wang et al. [63] focused on narrow or synthetic tasks for fine-tuning. Bricken et al. [4] studied the change in sleeper agent features via stage-wise model diffing. While these works focus on specific scenerios, there remain two limitations: (1) there lacks a general and unified way or pipeline for interpreting the learning process that is applicable to various scenarios from the view of mechanistic interpretability; (2) current studies mostly focus on post-hoc interpretations of fine-tuning, without daring to provide guidance for a more precise and effective fine-tuning process with interpretability. These limitations make it difficult for interpretability works to be well applied in practice to more general settings, or to provide more inspiration for traceable and steerable training.
To solve the above two concerns, we introduce a general and interpretable fine-tuning method, making it possible for transparent training. Specifically, inspired by prior works on intrinsic dimensionality, we view the learning process inside a model as finding and optimizing the minimum subgraph corresponding to the intrinsic dimensions in a computational graph. Previous estimation of intrinsic dimensionality was conducted directly in the parameter space, while the research on mechanistic interpretability usually adopts a representational perspective. To bridge these two paradigms, we introduce the node-level intrinsic dimensionality for the first time, enabling the study of learning dynamics with intrinsic dimensionality in the representation space.
Based on our theory, we propose circuit-tuning, a method that performs interpretable fine-tuning in a computational graph. Our method is able to: (1) precisely and dynamically localize the key parts in a computational graph during fine-tuning on any tasks and (2) construct and optimize the circuit to fit a new data distribution, thereby actively providing guidance for the fine-tuning process from an interpretable perspective. Experiments confirm the validity of our theory and the effectiveness of our method on various scenarios. We visualize and analyze a transparent, interpretable, and traceable training process, provide new findings about the self-organization mechanism inside a model during fine-tuning, and open new horizons for controllable and steerable training.
In summary, our contributions are mainly as follows:
  • We redefine intrinsic dimensionality and extend the estimation of it from the parameter space to the representation space, thereby introducing that concept into mechanistic interpretability and making it possible to study the learning dynamics in a computational graph.
  • We propose circuit-tuning, a general way of interpretable fine-tuning. Our method integrates heuristic and dynamic circuit discovery into fine-tuning for the study of learning dynamics for the first time, serving as a feasible way of interpretability-driven training.
  • We confirm the effectiveness of our method through experiments and gain new insights into the mechanisms behind learning through causal analysis and visualization, providing new insights and improvements for the study of learning dynamics.
2 Related work
2.1 Intrinsic dimensionality
Intrinsic dimension was studied in early works to estimate the dimensionality of a dataset [15, 58, 6, 29]. Inspired by the redundancy in parameters, Li et al. [31] defined intrinsic dimensionality as the minimum number of parameters needed to reach satisfactory solutions for an objective function. Aghajanyan et al. [1] focused on fine-tuning and defined it as the minimal description length in terms of a specific task, providing new insights into the generalization during pre-training. Prior works directly looked at the weights in a model and measure the intrinsic dimensionality in the parameter space, while the decomposition of a model and the following research on mechanistic interpretability focused on the activations in the representation space. The divergence of research perspectives has brought difficulties to the study of intrinsic dimensionality in the field of mechanistic interpretability.
2.2 Mechanistic interpretability for learning